OpenStudy (arihangdu):

Verify the identity.please show all works. Cot( theta - pai/2) = - tan theta.

OpenStudy (jchick):

Some identities: cos(a-b) = cos(a)cos(b) + sin(a)sin(b) sin(a-b) = sin(a)cos(b) - cos(a)sin(b) cot(a) = 1/tan(a) Therefore (substituting x for theta): cos(x-pi/2) = cos(x)cos(pi/2) + sin(x)sin(pi/2) ---------------------- sin(x-pi/2) = sin(x)cos(pi/2) - cos(x)sin(pi/2) cos(pi/2) =0 sin(pi/2) =1 So: sin(x)/-cos(x) = -tan(x)

OpenStudy (arihangdu):

Is this the way to do this?

OpenStudy (mathmale):

Yes. You need to learn trig identities such as those jchick has typed here for you, and know when and how to apply them. More questions?

OpenStudy (arihangdu):

I don't understand how comcos( X- pi/2)=cos(X)cos(pi/2)+sin(X)sin(pi/2) is

OpenStudy (arihangdu):

@jchik please can you explain me more about

OpenStudy (jchick):

For any CO - trigonometric functions m(x) and co-m(x), this property is always true. m(x)=co−m(π/2−x)=co−m(x−π/2).....or..vise..versa−−−−if..m(x)..is..even m(x)=co−m(π/2−x)=−co−m(x−π/2).....or..vise..versa−−−−if..m(x)..is..odd

OpenStudy (jchick):

cot ( theta - pi / 2 )= cos ( theta - pi / 2 ) / sin ( theta - pi / 2 ) cos ( theta - pi / 2 )= cos(theta) cos (pi/2) + sin(theta) sin (pi/2) = cos(theta) (0) + sin(theta) (1) = sin(theta) sin ( theta - pi / 2 )= sin(theta) cos(pi/2) - cos(theta) sin(pi/2) = sin(theta) (0) - cos(theta) (1) = -cos(theta) so you get, cot ( theta - pi / 2 )= sin(theta) / -cos(theta) cot ( theta - pi / 2 )= -tan(theta)

OpenStudy (jchick):

Do you get it now?

OpenStudy (arihangdu):

Thank you so much

OpenStudy (jchick):

No problem!

OpenStudy (arihangdu):

@ Nnesha too long you tap on May you doing other way as well?

Nnesha (nnesha):

there is another to prove this identity by using Cofunctions Identities where it says \[\large\rm \cot(\frac{\pi}{2}\color{Red}{-}\theta )=\tan x\] first rewrite the left side as \[\cot[\color{red}{-}(\frac{\pi}{2}-\theta)]\] i just take out the negative sign and then use the identity cot(pi/2-x)= tanx and use the fact cot is an odd function \[\large\rm \cot[\color{ReD}{-}(\frac{\pi}{2}-\theta)] \rightarrow \color{Red}{-} \cot(\frac{\pi}{2}-\theta)\]

Nnesha (nnesha):

~cot(-x)= -cot(x) ^^

OpenStudy (jchick):

Yes @Nnesha that would work as well.

OpenStudy (jchick):

I still need to work with using the equation tab. I am not used to having that function to teach with.

OpenStudy (arihangdu):

So, I confuse among this twos which is the most correct one?

OpenStudy (jchick):

Have you learned about the Confunctions?

OpenStudy (arihangdu):

Not yet

OpenStudy (jchick):

Ok then they probably aren't looking for it that way. So if you haven't learned about it put what I have as long as what I have is the way you have learned to do it.

OpenStudy (arihangdu):

Please show me I simple and understandable ways.

OpenStudy (jchick):

Do you understand what I have?

OpenStudy (arihangdu):

Of course I did

OpenStudy (jchick):

Ok then you have learned about it that way, only put what you know and if you know and understand what I have put that down.